Review 1 Constants | Periodic Table Part A James Cameron piloted a submersible craft to the bottom of the Challenger Deep, the deepest point on the oceans floor, 11,000 m below the surface. What was the total inward force on the 1.1-m-diameter pllot sphere in which Cameron sat? Express your answer with the appropriate units. F 2.84 10 N Submit Previous Answers Request Answer Incorrect; Try Again; 5 attempts remaining Provide Feedback
Answer #1

Solution)

h=11.000 m D=1.1m Pn = 1027 kg/m 3 The total surface area of the sphere is, = 4(3.14)| Lim 3.79 m2 The total pressure on the sphere is, (1.01x10 Pa)+(1027 kg/m3)(9.81 m/s)(11,000 m) -(1.01x 105 Pa)-(1108.2x10s Pa) 1.11×108 Pa The equation for relation of force, area and pressure is, F=PA Substitute the values in the above equation. F=PA =(1.11×108 Pa)(3.79 ㎡) =4.2x108 N Therefore, the total inward force on the sphere is 14.2×10° NI.